OpenStudy (samara8954):

There are 96 pencils packed in 8 boxes. How many boxes are needed for 180 pencils?

OpenStudy (aripotta):

make a proportion\[\frac{ pencils }{ boxes }\frac{ 96 }{ 8 }=\frac{ 180 }{ x }\]then cross-multiply and solve for x

OpenStudy (anonymous):

96 dived by 8 will tell you there is 12 pencils in each box

OpenStudy (aripotta):

proportions seem easier to me since it's also a visual to what's going on. but that's just me :)